# Combining Deterministic Enhanced Conditions with Dual-Streaming Encoding for Diffusion-Based Speech Enhancement

### v1

@@ IV-F Evaluation of Different Diffusion Models
     + Dual-Stream 27.3 25.2 3.65 2.85 2.98 12.01 2.34 0.954 3.69 3.44 4.09 3.18 3.77 3.60
IV-G Evaluation in the Proposed Deterministic Model
@@
Table II shows the performance of the proposed COFFEE method. We compared the effects of T-RF and T-F across different layers. Exp.–1 to Exp.–10 compared the impact of using T-RF in different decoder layers. From the first to the fifth decoder layer, simply stacking the network in different layers improved all evaluation metrics. Through Exp.–6 to Exp.–9, we compared the impact of T-RF in different decoder layers. The experimental results show that applying fine-grained processing closer to the output layer led to more significant performance improvements. Comparing Exp.–5 and Exp.–9, although Exp.–5 showed better performance, using fine-grained processing only in the final decoder layer provided advantages in terms of the number of parameters and processing speed, with minimal performance difference between these two models. Exp.–10 used a T-F processing block after the decoder. Compared to using T-RF blocks in every decoder layer (Exp.–5), using fine-grained processing only after the final decoder layer or after the decoder yielded better performance.
We compared the proposed method with other deterministic models in Table I. TF-GridNet is currently the best-performing model on the CHiME4 dataset. The proposed model had comparable performance to TF-GridNet. The proposed COFFEE outperformed DCCRN and DPRNN.
@@
Table III shows the performance of the proposed deterministic model for diffusion on the evaluation sets of the CHiME4 dataset. Compared to other deterministic models in Table I, the proposed COFFEE achieved better dMOS and uMOS values on the real evaluation sets. On the real evaluation sets, Deter-Diffusion with the proposed COFFEE achieved better dMOS values compared to other baseline deterministic models. However, it only outperformed DCCRN for uMOS. Compared to other baseline deterministic models, Deter-Diffusion with COFFEE had better noise suppression (Cbak). On the simulated evaluation sets, Deter-Diffusion with COFFEE achieved the best performance compared to other deterministic models. For StoRM and GP-Unified, using COFFEE as the deterministic model achieved better performance on both simulated and real evaluation sets compared to using the same diffusion model with other deterministic models. In Table I, TF-GridNet with fine-grained processing did not achieve the best performance, while diffusion models with DCCRN demonstrated more stable performance, which motivated the design of COFFEE. Combining fine-grained and coarse-grained processing demonstrated strong performance: diffusion models with COFFEE outperformed those diffusion models using other deterministic models. In uMOS evaluations, diffusion models with COFFEE had the same trend as other deterministic models: RDM-SE $>$ GP-Unified $>$ StoRM.
@@ V Conclusions
Conventional diffusion models need to incorporate accurate prior knowledge as reliable conditions to generate accurate predictions. In this paper, we propose the Repair-Diffusion Model for Speech Enhancement (RDM-SE) to provide reliable conditions for diffusion using the deterministic model. It explicitly repairs deterministic information using an additional deterministic decoder. We first experimented with using DCCRN, DPRNN, and TF-GridNet as deterministic models to evaluate their impact on diffusion models on CHiME4 datasets. We found that while using diffusion models did not necessarily enhance human hearing experiences on simulated evaluation sets, it had certain advantages on real evaluation sets. The proposed RDM-SE was compared with Deter-Diffusion, which directly uses enhanced deterministic features as conditions, as well as with StoRM and GP-Unified, which use both noisy and enhanced deterministic features as conditions. The proposed RDM-SE had better noise suppression and further improved both dMOS and uMOS values compared to StoRM and GP-Unified. Deterministic models had a significant impact on diffusion performance The experimental results also suggested that the choice between deterministic-only and deterministic-noisy conditions depends on the specific deterministic model. Thus, we further improved the RDM-SE with dual-stream encoding to utilize both deterministic-only and deterministic-noisy conditions (DERDM-SE). Furthermore, deterministic models with fine-grained processing had unstable performance on real evaluation sets, whereas coarse-grained processing based on the UNet structure had more stable diffusion performance. Therefore, we proposed a deterministic model that combines coarse-grained and fine-grained processing, called COarse Fisrt then Fine EnhancEment (COFFEE). With COFFEE as the deterministic model, different diffusion models had stable and superior performance compared to those using other deterministic models.

### v2

@@ IV-C Settings for the Proposed Methods
   + D3U-SE 27.3 25.2 3.65 2.85 2.98 12.01 2.34 0.954 3.69 3.44 4.09 3.18 3.77 3.60
IV-D Additional Settings for Evaluation
In the diffusion process, we used PC sampling with 30 steps. In previous work [18], we found that within a 30-step reverse diffusion process, the sampling diversity at step 20 (with $t$ progressing from $T$ to 0) can be maintained as if starting from the initial state. Thus, in this paper, we start the diffusion process at $t=20$ instead of $t=30$. In previous work [18], we found that when the ensemble number was greater than 8, the performance difference became negligible. Therefore, this paper used the ensemble with a maximum of 8 samples.
We test for statistical significance using a two-sided paired $t$-test with $n=330$ (real) and $n=420$ (simulated) for uMOS/dMOS. With sample standard deviation of paired differences $s_{\Delta}=0.50$, the standard error is $\mathrm{SE}=s_{\Delta}/\sqrt{n}\approx 0.0275$ (real) and $0.0244$ (simulated). Hence an average gain $\Delta\mathrm{MOS}\geq 0.10$ is statistically significant (its 95% confidence interval excludes 0) for both sets.
IV-E Evaluation Metrics
Multiple evaluation metrics were used to analyze the improvements in human listening experience. For the simulated data, signal distortion (Csig) [39], background intrusiveness (Cbak) [39], overall quality (Covl) [39], the wideband perceptual evaluation of speech quality (W-P) [40], the short-time objective intelligibility (STOI) [41], and signal to distortion ratio (SDR) were used as evaluation metrics. For the real data, the Mean Opinion Score was used, which was evaluated with the open-source toolkit DNSMOS (dMOS) [42, 43]. UTMOS (uMOS) [44] was also used. Although both the dMOS and uMOS are designed to approximate subjective perception, they still remain objective metrics.
IV-F Evaluation in the Proposed Deterministic Model
@@
Table I shows the performance of the proposed COFFEE method. We compared the effects of T-RF and T-F across different layers. Exp.–1 to Exp.–10 compared the impact of using T-RF in different decoder layers. From the first to the fifth decoder layer, simply stacking the network in different layers improved all evaluation metrics. Through Exp.–6 to Exp.–9, we compared the impact of T-RF in different decoder layers. The experimental results show that applying fine-grained processing closer to the output layer led to more significant performance improvements. Comparing Exp.–5 and Exp.–9, although Exp.–5 showed better performance, using fine-grained processing only in the final decoder layer provided advantages in terms of the number of parameters and processing speed, with minimal performance difference between these two models. Exp.–10 used a T-F processing block after the decoder. Compared to using T-RF blocks in every decoder layer (Exp.–5), using fine-grained processing only after the final decoder layer or after the decoder yielded better performance. We compared the proposed method with other deterministic models. TF-GridNet is currently the best-performing model on the CHiME4 dataset. The proposed model had comparable performance to TF-GridNet. The proposed COFFEE outperformed DCCRN and DPRNN.
IV-G Evaluation of Different Diffusion Models
Table II shows the performance of different deterministic models in combination with various diffusion models. Compared to the noisy mixture signal, SGMSE+ effectively enhances noisy signals. Deter-Diffusion, which used only deterministic enhanced features as conditions, performed better than SGMSE+. When using DCCRN-enhanced features as conditions, the performance of objective evaluation metrics (Csig, Cbak, Covl, SDR, WB-PESQ, STOI) on the simulated dataset was comparable to that of SGMSE+. This may be due to the signal distortions introduced by DCCRN. The DCCRN-Deter-Diffusion performed better at noise reduction, achieving the higher Cbak value, while the noisy SGMSE+ method excelled at speech signal recovery, yielding the higher Csig value. Although DCCRN-Deter-Diffusion did not outperform SGMSE+ in simulated objective evaluation metrics, it showed certain advantages in MOS: Deter-Diffusion achieved higher scores on both dMOS and uMOS on the real evaluation sets. The quality of the deterministic enhanced features affected the final performance of Deter-Diffusion. The performance ranking of the deterministic models on the simulated evaluation sets: TF-GridNet >> DPRNN >> DCCRN. However, TF-GridNet-Deter-Diffusion performed the worst on the simulated data. While the performance on the real evaluation sets was inconsistent with that on the simulated evaluation sets: DPRNN >> TF-GridNet >> DCCRN. TF-GridNet outperformed DPRNN in dMOS but underperformed in uMOS. The TF-GridNet-Deter-Diffusion and DPRNN-Deter-Diffusion performed worse in dMOS on the real evaluation sets than the DCCRN-Deter-Diffusion.
We also compared two models that incorporate the original noisy features as conditions: StoRM and GP-Unified. The models with deterministic-noisy conditions demonstrated greater sensitivity to the deterministic enhanced features, regardless of whether deterministic information was incorporated into the diffusion model implicitly or explicitly during training.
GP-Unified explicitly incorporates deterministic information by using a deterministic decoder during training. Compared to StoRM, GP-Unified had comparable performance on the simulated evaluation sets but had better performance on the real evaluation sets. For all dMOS and uMOS values on the real evaluation sets, GP-Unified outperformed StoRM. However, compared to Deter-Diffusion, GP-Unified still showed a decline in some metrics. Only with TF-GridNet-enhanced conditions, GP-Unified achieved better performance on the simulated evaluation sets (WB-PESQ, STOI). For the real evaluation sets, GP-Unified still had some disadvantages in dMOS values, which were similar to those of StoRM: with DCCRN or DPRNN conditions, both StoRM and GP-Unified had a decline in dMOS values. By comparing Deter-Diffusion, StoRM, and GP-Unified across various metrics on the real evaluation sets, we found significant declines in noise reduction (Cbak) for StoRM and GP-Unified. This may explain the decrease in dMOS values: [45] indicates that the uMOS is insensitive to speech with noise compared to dMOS. For uMOS values, with DCCRN or DPRNN conditions, GP-Unified outperformed Deter-Diffusion. However, with TF-GridNet conditions, although it showed some improvement over StoRM, it still had significant performance degradation compared to Deter-Diffusion. Based on the performance of TF-GridNet as the condition on simulated and real evaluation sets, we argue that the T-F bin-level fine-grained processing, while achieving best performance on simulated data, may introduce distortions more easily to real data due to distribution differences between real and simulated data. However, the advantages brought by fine-level processing cannot be ignored for deterministic models. Besides, although DCCRN has lower dMOS and uMOS values than TF-GridNet in real evaluation sets, its performance was acceptable and demonstrates more stable diffusion performance, particularly in models that combine deterministic-noisy features.
(a) dMOS score on BUS noise scenario
(b) dMOS score on CAF noise scenario
(c) dMOS score on PED noise scenario
(d) dMOS score on STR noise scenario
Figure 4: dMOS scores under different noise scenarios using various SE systems on the real evaluation set of the CHiME4 dataset. In each group, models below the red dashed line exhibit a statistically significant performance improvement (significance computed as described in Section IV-D). Bars marked with denote the proposed model.
(a) uMOS score on BUS noise scenario
(b) uMOS score on CAF noise scenario
(c) uMOS score on PED noise scenario
(d) uMOS score on STR noise scenario
Figure 5: uMOS scores under different noise scenarios using various SE systems on the real evaluation set of the CHiME4 dataset. In each group, models below the red dashed line exhibit a statistically significant performance improvement (significance computed as described in Section IV-D). Bars marked with denote the proposed model.
Table II also shows our proposed D2U-SE. We also used DCCRN, DPRNN, and TF-GridNet as deterministic models for diffusion. Compared to StoRM and GP-Unified, the proposed D2U-SE achieved better and more stable performance. On the real evaluation sets, D2U-SE outperformed StoRM and GP-Unified in almost all deterministic models. Compared to StoRM and GP-Unified, D2U-SE achieved better noise suppression, particularly on the real evaluation sets, with higher Cbak values. D2U-SE showed a greater advantage in uMOS compared to StoRM and GP-Unified: using all deterministic models, D2U-SE achieved consistent improvements compared to other deterministic-noisy diffusion models. For dMOS, different deterministic models had varying impacts on D2U-SE: with DCCRN and DPRNN, D2U-SE outperformed StoRM and GP-Unified; while with TF-GridNet, D2U-SE had comparable performance to StoRM and GP-Unified. Compared to Deter-Diffusion, D2U-SE showed significant improvements when using DCCRN, comparable performance when using DPRNN, and a slight decline when using TF-GridNet. Surprisingly, D2U-SE with DCCRN outperformed those using DPRNN and TF-GridNet, achieving the best performance. This did not occur in StoRM or GP-Unified. As previously analyzed, fine-grained processing may result in more severe distortion issues when handling real data. This suggests that while the proposed D2U-SE can mitigate distortion issues to some extent, its performance is still closely tied to the input deterministic feature. Explicitly incorporating deterministic repair ($\text{D}_{c}$ Decoder) further improved the model’s performance, especially when TF-GridNet was used as the condition. In addition, given the performance differences among different deterministic methods under deterministic-only conditions (Deter-Diffusion) and deterministic-noisy conditions (StoRM and GP-Unified), we proposed to use dual-stream encoding (D3U-SE). Although the dual-stream encoding model showed some performance degradation on simulated metrics (Csig, Cbak, Covl, SDR, WB-PESQ, STOI) compared to D2U-SE ($\text{D}_{c}$ Decoder), it achieved improvements in dMOS and uMOS.
TABLE III: The summary performance with Mean ($\pm$ Standard Deviation) of different methods under evaluation sets.
Systems Simulated Real
dMOS dMOS uMOS
Mixture 2.69 ($\pm$0.18) 2.47 ($\pm$0.22) 1.58 ($\pm$0.51)
SGMSE+ 3.54 ($\pm$0.23) 3.45 ($\pm$0.31) 2.63 ($\pm$0.65)
DCCRN 3.26 ($\pm$0.22) 3.25 ($\pm$0.30) 2.59 ($\pm$0.61)
   + Deter-Diffusion 3.63 ($\pm$0.20) 3.66 ($\pm$0.30) 2.93 ($\pm$0.59)
   + StoRM 3.60 ($\pm$0.21) 3.57 ($\pm$0.29) 3.16 ($\pm$0.50)
   + GP-Unified 3.58 ($\pm$0.20) 3.58 ($\pm$0.26) 3.28 ($\pm$0.45)
   + D3U-SE 3.66 ($\pm$0.21) 3.73 ($\pm$0.26) 3.49 ($\pm$0.49)
DPRNN 3.34 ($\pm$0.20) 3.07 ($\pm$0.23) 2.90 ($\pm$0.54)
   + Deter-Diffusion 3.65 ($\pm$0.21) 3.65 ($\pm$0.25) 3.36 ($\pm$0.48)
   + StoRM 3.61 ($\pm$0.21) 3.56 ($\pm$0.25) 3.32 ($\pm$0.48)
   + GP-Unified 3.66 ($\pm$0.21) 3.61 ($\pm$0.23) 3.42 ($\pm$0.45)
   + D3U-SE 3.66 ($\pm$0.21) 3.69 ($\pm$0.25) 3.53 ($\pm$0.48)
TF-GridNet 3.55 ($\pm$0.23) 3.35 ($\pm$0.43) 2.71 ($\pm$0.85)
   + Deter-Diffusion 3.60 ($\pm$0.20) 3.56 ($\pm$0.23) 3.34 ($\pm$0.48)
   + StoRM 3.67 ($\pm$0.20) 3.52 ($\pm$0.39) 2.89 ($\pm$0.75)
   + GP-Unified 3.65 ($\pm$0.21) 3.52 ($\pm$0.39) 2.96 ($\pm$0.71)
   + D3U-SE 3.70 ($\pm$0.21) 3.58 ($\pm$0.39) 3.09 ($\pm$0.75)
COFFEE 3.43 ($\pm$0.21) 3.47 ($\pm$0.29) 3.20 ($\pm$0.48)
   + Deter-Diffusion 3.68 ($\pm$0.19) 3.75 ($\pm$0.26) 3.23 ($\pm$0.55)
   + StoRM 3.72 ($\pm$0.20) 3.80 ($\pm$0.24) 3.43 ($\pm$0.47)
   + GP-Unified 3.70 ($\pm$0.20) 3.72 ($\pm$0.26) 3.47 ($\pm$0.44)
   + D3U-SE 3.69 ($\pm$0.20) 3.77 ($\pm$0.25) 3.60 ($\pm$0.46)
Figure 6: Heatmap showing the differences in Csig, Cbak and Covl scores when evaluated using DNSMOS and the Composite toolkit (requiring clean labels) under simulated evaluation sets. The difference values are calculated as (DNSMOS estimated score - Composite toolkit estimated score).
@@ IV-G Evaluation of Different Diffusion Models
Table II shows the performance of the proposed deterministic model for diffusion on the evaluation sets of the CHiME4 dataset. Compared to other deterministic models, the proposed COFFEE achieved better dMOS and uMOS values on the real evaluation sets. On the real evaluation sets, Deter-Diffusion with the proposed COFFEE achieved better dMOS values compared to other baseline deterministic models. However, it only outperformed DCCRN for uMOS. Compared to other baseline deterministic models, Deter-Diffusion with COFFEE had better noise suppression (Cbak). On the simulated evaluation sets, Deter-Diffusion with COFFEE achieved the best performance compared to other deterministic models. For StoRM and GP-Unified, using COFFEE as the deterministic model achieved better performance on both simulated and real evaluation sets compared to using the same diffusion model with other deterministic models. TF-GridNet with fine-grained processing did not achieve the best performance, while diffusion models with DCCRN demonstrated more stable performance, which motivated the design of COFFEE. Combining fine-grained and coarse-grained processing demonstrated strong performance: diffusion models with COFFEE outperformed those diffusion models using other deterministic models. In uMOS evaluations, diffusion models with COFFEE had the same trend as other deterministic models: D2U-SE >> GP-Unified >> StoRM. In addition, we also evaluated the standard deviations of different methods under both dMOS and uMOS, which is shown in Table III. The results show that, except when TF-GridNet was used as the condition—where the proposed D3U-SE experiences a performance drop compared to Deter-Diffusion—the performance improves while maintaining stability under other deterministic conditions, especially in terms of the uMOS metric.
Fig. 4 and Fig. 5 compare the dMOS and uMOS performance of different SE models under various noise scenarios. Among all noise conditions, the BUS noise scenario was consistently the most challenging scenario for all models to recover effectively. In addition, dMOS appears to be more difficult to improve compared to uMOS: under the uMOS metric, the proposed D2U-SE and D3U-SE models showed statistically significant improvements in most cases when conditioned on different deterministic models. However, under the dMOS metric, significant improvements are observed only when using DCCRN or DPRNN as the conditioning models. COFFEE demonstrates greater stability compared to other deterministic models: when used as the conditioning model, the dMOS scores across different diffusion models show no significant variation. When conditioned on TF-GridNet, the system showed greater instability in the CAF noise scenario: using the deterministic-only input significantly outperforms the deterministic-noisy condition. In contrast, this performance gap is much less pronounced in other noise environments.
IV-H Comparison of DNSMOS and Composite Toolkit Metrics
In Table II, the Csig, Cbak, and Covl values are computed using the Composite toolkit with the clean labels. The difference heatmap (DNSMOS estimated score - Composite toolkit estimated score) across the Csig, Cbak, and Covl is shown in Fig. 6. For all enhancement systems, Cbak shows a positive gain, indicating that DNSMOS tends to underestimate the amount of noise in the enhanced signals. The overall quality score Covl shows a substantial overestimation of overall speech quality by DNSMOS. In contrast, the speech signal quality score Csig is negative for most systems, indicating that DNSMOS underestimates speech quality. An interesting observation is that deterministic models generally yield smaller Cbak and Covl differences than diffusion models that use them as the corresponding condition, but larger differences in Csig. This observation suggests that, to achieve more stable performance—particularly in certain downstream tasks—it may be necessary to introduce a bias or set higher or lower thresholds for evaluation metrics like DNSMOS and UTMOS, relative to conventional objective measures.
@@ V Conclusions
Conventional diffusion models need to incorporate accurate prior knowledge as reliable conditions to generate accurate predictions. In this paper, we propose a Deterministic-Diffusion Unified model for Speech Enhancement (D2U-SE) to provide reliable conditions for diffusion using the deterministic model. It explicitly incorporates deterministic information using an additional deterministic decoder. We first experimented with using DCCRN, DPRNN, and TF-GridNet as deterministic models to evaluate their impact on diffusion models on CHiME4 datasets. We found that while using diffusion models did not necessarily enhance objective metrics on simulated evaluation sets, it had certain advantages on real evaluation sets. The proposed D2U-SE was compared with Deter-Diffusion, which directly uses enhanced deterministic features as conditions, as well as with StoRM and GP-Unified, which use both noisy and enhanced deterministic features as conditions. The proposed D2U-SE had better noise suppression and further improved both dMOS and uMOS values compared to StoRM and GP-Unified. Deterministic models had a significant impact on diffusion performance. The experimental results also suggested that the choice between deterministic-only and deterministic-noisy conditions depends on the specific deterministic model. Thus, we further improved the D2U-SE with dual-stream encoding to utilize both deterministic-only and deterministic-noisy conditions (D3U-SE). Furthermore, deterministic models with fine-grained processing had unstable performance on real evaluation sets, whereas coarse-grained processing based on the UNet structure had more stable diffusion performance. Therefore, we proposed a deterministic model that combines coarse-grained and fine-grained processing, called COarse First then Fine EnhancEment (COFFEE). With COFFEE as the deterministic model, different diffusion models had stable and superior performance compared to those using other deterministic models.